# WMNav: Integrating Vision-Language Models into World Models for Object Goal Navigation

### v4

@@ I Introduction
The world model is a computational representation of environment dynamics that equips agents with the ability to simulate interactions. By simulating action choices within the virtual environment, agents can explore potential outcomes safely without directly interacting with the environment. This method not only reduces collision with obstacles but also enhances the agent’s efficiency in planning and exploring. However, the true challenge lies in creating a versatile world model that can faithfully capture the landscape of an indoor environment. blocksworld [16] and gridworld [17] prove that LLMs can function as effective world models in simplistic settings. WebDreamer [18] uses LLMs as world models in web environments for model-based planning of web agents, demonstrating its effectiveness over other baselines. Can VLMs also function as world models? Given that VLMs are trained on vast amounts of first-person human perspective images, we hypothesize that they have acquired sufficient common sense knowledge to simulate the outcomes resulting from human choices and actions. With their extensive pre-trained knowledge spanning indoor layout and general reasoning ability, VLM has great potential to take on this task.
@@ II-B Foundation Model Guided Navigation
Recent advances in embodied navigation highlight the complementary roles of Vision-Language Models (VLMs) and Large Language Models (LLMs), driven by their distinct strengths in visual grounding and common sense reasoning. LLMs enable zero-shot commonsense reasoning through object-room correlation prediction [5], semantic mapping [28, 7], and chain-of-thought planning [29, 30]. However, LLMs are unable to process rich visual information and do spatial reasoning. VLMs like CLIP [22, 2, 31] address this by directly aligning visual observations with textual goals through multimodal embeddings, offering seamless integration with exploration policies through unified visual features. Compared to LLM-based methods, VLM-based methods [9, 13] may be more efficient and elegant for target-driven navigation by avoiding language-mediated grounding gaps. This is due to their multimodal nature and better spatial reasoning capabilities. Our framework uses pure VLM, and unlike other approaches, we use the VLM to predict the state of the environment caused by actions.
@@ III-C2 Curiosity Value Map Construction
The construction and update process is illustrated in Figure 3 (c). Initially, all pixels in $M^{cv}_{0}$ are set to 10, indicating it is possible to find the target in all regions as the agent does not have any information about the entire scene. To convert the scores from the ego-centric perspective to the top-down view of the curiosity value map, we perform a projection transformation using depth information $D_{t}$ and pose $P_{t}$, similar to [13]. We project the navigable area with scores $Score_{t}$ from the ego-centric view onto a top-down map $M^{nav}_{t}$ as in Figure 3 (b):
@@ IV-D Ablation Study
Effect of different VLMs. We further evaluate the abilities of different VLMs in navigation as shown in Table III. For each row of the table, all the VLMs used in the framework are guaranteed to be of the same model. Gemini 1.5 Pro (row 3) demonstrates superior capability in the task. Note that when using the smaller Gemini 1.5 Flash VLM, our approach still achieves competitive performance compared to other methods on HM3D, which means our framework is practical on its own rather than relying solely on the capabilities of VLM. What’s more, With the evolution of open-source and proprietary VLMs, the ability of each module of our model can be significantly enhanced, and as a whole, it has the potential to show better performance.

### v5

@@ I Introduction
The world model is a computational representation of environment dynamics that equips agents with the ability to simulate interactions. By simulating action choices within the virtual environment, agents can explore potential outcomes safely without directly interacting with the environment. This method not only reduces collision with obstacles but also enhances the agent’s efficiency in planning and exploring. However, the true challenge lies in creating a versatile world model that can faithfully capture the landscape of an indoor environment. ATD [16] uses a dual-branch self-guided imagination policy based on large language models with a human-like left-right brain architecture. WebDreamer [17] uses LLMs as world models in web environments for model-based planning of web agents, demonstrating its effectiveness over other baselines. Can VLMs also function as world models? Given that VLMs are trained on vast amounts of first-person human perspective images, we hypothesize that they have acquired sufficient common sense knowledge to simulate the outcomes resulting from human choices and actions. With their extensive pre-trained knowledge spanning indoor layout and general reasoning ability, VLM has great potential to take on this task.
@@ II-B Foundation Model Guided Navigation
Recent advances in embodied navigation highlight the complementary roles of Vision-Language Models (VLMs) and Large Language Models (LLMs), driven by their distinct strengths in visual grounding and common sense reasoning. LLMs enable zero-shot commonsense reasoning through object-room correlation prediction [5], semantic mapping [27, 7], and chain-of-thought planning [28, 29]. However, LLMs are unable to process rich visual information and do spatial reasoning. VLMs like CLIP [21, 2, 30] address this by directly aligning visual observations with textual goals through multimodal embeddings, offering seamless integration with exploration policies through unified visual features. Compared to LLM-based methods, VLM-based methods [9, 13] may be more efficient and elegant for target-driven navigation by avoiding language-mediated grounding gaps. This is due to their multimodal nature and better spatial reasoning capabilities [31]. Our framework uses pure VLM, and unlike other approaches, we use the VLM to predict the state of the environment caused by actions.
@@ III-C2 Curiosity Value Map Construction
The construction and update process is illustrated in Figure 3 (c). Initially, all pixels in $M^{cv}_{0}$ are set to 10, indicating it is possible to find the target in all regions as the agent does not have any information about the entire scene. To convert the scores from the ego-centric perspective to the top-down view of the curiosity value map, we perform a projection transformation using depth information $D_{t}$ and pose $P_{t}$, similar to [13] [37]. We project the navigable area with scores $Score_{t}$ from the ego-centric view onto a top-down map $M^{nav}_{t}$ as in Figure 3 (b):
@@ IV-D Ablation Study
Effect of different VLMs. We further evaluate the abilities of different VLMs in navigation as shown in Table III, including open-source models and proprietary models. For each row of the table, all the VLMs used in the framework are guaranteed to be of the same model. The comparison between the 3B and 7B models of Qwen2.5-VL [43] shows that an increase in the model’s scale can effectively enhance the agent’s capabilities. Comparison between Qwen2.5-VL and the Gemini series models reveals that current open-source multimodal models still lag behind advanced proprietary models in terms of ZSON. Gemini 1.5 Pro demonstrates superior capability in the task. Note that when using the smaller Gemini 1.5 Flash VLM, our approach still achieves competitive performance compared to other methods on HM3D, which means our framework is practical on its own rather than relying solely on the capabilities of VLM. What’s more, With the evolution of open-source and proprietary VLMs, the ability of each module of our model can be significantly enhanced, and as a whole, it has the potential to show better performance.